Design of Algorithms - Course Review

K. Subramani

LCSEE, West Virginia University,

Morgantown, WV {ksmani@csee.wvu.edu}

February 10, 2011

1. Are you comfortable with the pace of the class? (a) Yes.

(b) No.

(c) Most of the time.

2. If you are not comfortable with the pace, have you interrupted the instructor in class?

(a) Yes, I have.

(b) No, it is entirely the instructor’s fault if he

cannot read my mind.

3. Is the course challenging and interesting?

(a) Yes.

(b) No.

(c) Sometimes.

(d) Most of the time.

4. Does the instructor address your concerns in timely fashion?

(a) Yes.

(b) No.

(c) Most of the time.

5. Is the instructor helpful?

(a) Yes.

(b) No.

(c) Most of the time.

(d) Sometimes.

6. How much time do you spend on this subject outside of class?

(a) None.

(b) 1-2 hours per week.

(c) 4-5 hours per week.

7. Are you learning new material in this course? (a) A lot.

(b) Some.

(c) No.

8. Are your critical thinking skills improving through this course?

(a) Yes, significantly.

(b) Somewhat.

(c) No.

9. Do you work on problems presented in class? (a) Yes.

(b) No, I prefer to solve them in my head.

10. Have you missed more than three classes?

(a) Yes.

(b) No.

11. Do you arrive on time for the class?

(a) Always.

(b) Most of the time.

(c) Rarely.

(d) Never.

12. Do you carefully review the material covered in previous classes before attending class?

(a) Never.

(b) Rarely.

(c) Most of the time.

(d) Always.

13. Have the pre-requsite courses prepared you properly for this class?

(a) No, not even close. The pre-reqs need to be strengthened.

(b) For the most part.

(c) Yes.

14. Is the lecturer effective at communicating ideas? (a) Always.

(b) Most of the time.

(c) Sometimes.

(d) Never.

15. Is the lecturer well-organized?

(a) Always.

(b) Never.

(c) Most of the time.

(d) Sometimes.

16. Have you met the instructor during office hours? (a) Occasionally.

(b) Never.

17. How would you rate this instructor vis-a-vis other instructors that you have had?

(a) Better than other instructors.

(b) As good as other instructors.

(c) Worse than other instructors.

18. How would you rate this course vis-a-vis other courses that you have taken?

(a) Better than other courses.

(b) As good as other courses.

(c) Worse than other courses.

19. Does the instructor show mastery over the material?

(a) Yes.

(b) No.

20. Is the instructor enthusiastic about the material? (a) Always.

(b) Most of the time.

(c) Sometimes.

(d) Never.

21. Would you take another course from this instructor?

(a) Definitely.

(b) Highly likely.

(c) Unlikely.

(d) Only if hell freezes over.

22. What aspects of the course would you like to change and in what way?